OpenStudy (anonymous):

The drama teacher is making costumes for the school play. Each costume requires 1 3/4 yards of material and 2/3 of a yard of ribbon.. The material costs 2.99 per yard and the ribbon costs .59 per yard. What is the approximate cost of each costume? i have to show my work

OpenStudy (texaschic101):

turn your fractions into decimals 1 3/4 = 1.75 2/3 = 0.67 now multiply .. 1.75(2.99) + 0.67(0.59) = T 5.23 + .40 = $5.63

OpenStudy (anonymous):

thank u very much